Fill in the blanks.
(a) (–8) + _____ = 0
(b) 13 + _____ = 0
(c) 12 + (–12) = ____
 
 
iaemfuewgnuersiojmhnbiorjhmneru

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(a) (-8) + 8 = 0

(b) 13 + (-13) = 0

(c) 12 + (-12) = 0



​Thanks